OpenStudy (anonymous):

25^2c = sqrt5 ^(4c + 16)

Parth (parthkohli):

Try to get the same base first.

OpenStudy (lgbasallote):

25 = 5^2 \(\sqrt 5\) = 5^1/2 so just rewrite

Parth (parthkohli):

\(\color{Black}{\Rightarrow (5^2)^{2c} =(5^{0.5})^{4c + 16} }\)

OpenStudy (lgbasallote):

\[25^{2c} = 5^{2(2c)}\] \[\sqrt5^{(4c + 16)} = 5^{1/2(4c + 16)}\]

Parth (parthkohli):

Now use the property: \(\color{Black}{\Rightarrow \large (a^x)^y = a^{xy} }\)

OpenStudy (lgbasallote):

then use the property \[a^x = a^y \implies x = y\]

OpenStudy (anonymous):

you can also eliminate the base, since they're the same. and just solve for c ^^

Parth (parthkohli):

If the base is same, and we have an equals sign in between, it is obv that we'll have the same exponent.

OpenStudy (anonymous):

wow... tweeeenz!

OpenStudy (anonymous):

How do I change it to an improper fraction? I haven't done it in two years.

OpenStudy (lgbasallote):

improper fraction?

OpenStudy (lgbasallote):

where will you use that?

OpenStudy (anonymous):

My answer choices are: a. c = 4 b. c = 7/2 c. c = -7/2 d. c = -5/2

OpenStudy (anonymous):

well it depends. what was your answer for c?

OpenStudy (anonymous):

What's that supposed to mean?

OpenStudy (anonymous):

well. okay in all cases. if you have an answer like: 2 and a quarter. you have 2 and 1/4. so you multiply the 2 (the whole number) with the 4 (the denominator), and then add whatever's on the numerator. so the improper fraction of 2 and 1/4 is 9/4.